We consider the following perturbations of a domain: bending a segment on the boundary; introducing a small slit in the interior; and stretching the domain along an axis. We consider analyticity and convergence of the associated Neumann/Dirichlet eigenvalue problems. As an application we show that almost every n-gon has simple Laplace spectrum. (Joint work with Luc Hillairet.)